OpenStudy (anonymous):

A chicken manages to prove that: \[ \sum_{i=1}^{n}0 = 0 \]where \(n\in \mathbb{N}\). She then continues by saying: \[ \lim_{n\to \infty}\sum_{i=1}^{n}0 = 0 \]In other words: \[ \underbrace{0+0+\dots+0}_{\infty} = 0 \]Is Mrs. Chicken correct? If not, what is the fallacy?

OpenStudy (anonymous):

@anyone_with_a_brain

OpenStudy (anonymous):

Woah , An SS 92 with a question

OpenStudy (anonymous):

This is the hardest math I have ever seen. What math is this?

OpenStudy (anonymous):

it's summation

OpenStudy (anonymous):

Her proof is as follows: \[ \sum_1^10=0 \]So \(p(1)\) is true. \[ \sum_1^{n+1}0=0+\sum_1^{n}0=0 \]So \(p(k)\implies p(k+1)\).

OpenStudy (anonymous):

This question is like something satellite would ask :P

OpenStudy (anonymous):

this question is like the key for immortality

OpenStudy (anonymous):

ALL right all SS 90 below , stop bombarding this question with unnecessary comments , let the pros do their work :D

OpenStudy (anonymous):

I already know the answer, it's just an interesting fact that is often overlooked.

OpenStudy (anonymous):

@wio , are u in a university or at school ?

hartnn (hartnn):

when you say lim n->infinity, you r just considering very large values of n so, when very large numbers of 0's are added, it has to give infinity

hartnn (hartnn):

has to give 0, i mean***

OpenStudy (anonymous):

Do you agree with the chicken?

OpenStudy (anonymous):

hartnn has a point , no matte how many times u add zero , it would still result in zero , Guess the chicken has brains

hartnn (hartnn):

yeah...

OpenStudy (anonymous):

What if the chicken says: \[ \underbrace{0+0+\dots+0}_\infty=\infty \times 0=0 \]

OpenStudy (anonymous):

isn't that what u already said the chicken said?

hartnn (hartnn):

that would be incorrect

OpenStudy (anonymous):

LOL? Of course 0+0+0+0+0+0+0=0... 0 has no value..

OpenStudy (anonymous):

@JOELMATHEWS1234 It's an indeterminate form.

hartnn (hartnn):

infinity is just a concept, there's no arithmetic with it

OpenStudy (anonymous):

So @hartnn , why would that be incorrect , is there some rule with indeterminates

OpenStudy (anonymous):

You can do arithmetic with infinity.

OpenStudy (anonymous):

Only with multiplying it to 0.

OpenStudy (anonymous):

or dividing it against itself to equal 1.

OpenStudy (anonymous):

@wio & @hartnn , according to tan graph , 1/0 = infinity , so is 0 * infinity = 1 ???

OpenStudy (anonymous):

Just because \(0/0\) is an indeterminate form doesn't mean you can't do arithmetic with integers.

OpenStudy (anonymous):

@ⒶArchie☁✪

OpenStudy (anonymous):

So you object to: \[ \underbrace{0+0+\dots+0}_\infty=\infty \times 0 \]@hartnn

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

what i was saying is just , infinity is alias for 'very large number' so, in that sense, when we add a very large numbers of 0, (we should get answer = 0) thats left side on right side, the notion is very large number times 0, which is also = 0 so in that sense i would agree to that

OpenStudy (anonymous):

\[ \sum_{i=0}^n\frac 1n=1 \]\(\lim_{n\to\infty}\) Gives: \[ \underbrace{\frac 1\infty+\frac 1\infty+\dots+\frac 1\infty}_{\infty}=1 \]

OpenStudy (anonymous):

Do you think it is impossible to add an infinite number of \(0\)s and eventually get something \(\neq 0\)?

OpenStudy (yttrium):

@wio, I think that will give us infinity, not 1.

OpenStudy (anonymous):

Nope, it gives \(n\times \frac 1n=1\)

hartnn (hartnn):

thats sum is not 1

OpenStudy (anonymous):

Make it \(i=1\)

OpenStudy (anonymous):

If you have \(i=0\) it is \(1+\frac 1n\) but I mean to say \(i=1\).

hartnn (hartnn):

still it has to be infinity, its diverging

OpenStudy (anonymous):

Can you prove it is diverging?

OpenStudy (anonymous):

Are you thinking of :\[ \sum_{i=1}^n\frac 1i \]

hartnn (hartnn):

oops, yes

OpenStudy (anonymous):

Okay so the chicken is correct that in this case \(\lim_{n\to\infty}n\times 0 = 0\)

hartnn (hartnn):

yes, the limiting value is indeed 0

OpenStudy (anonymous):

But is it true that \[ \sum_{i=1}^nc=n\times c \]Even when \(n\to \infty\)?

OpenStudy (anonymous):

@Yttrium , the sum is 1 , you're adding 1 infinite times, and then dividing by infinite , so it is in fact 1

hartnn (hartnn):

when n->infinity, it'll be c times infinity= infinity

OpenStudy (anonymous):

Except when \(c\leq0\)? Are you sure though?

hartnn (hartnn):

yeah, for negative c, it'll be negative infinity

OpenStudy (anonymous):

Hmm, I'm wondering how one would prove it.

hartnn (hartnn):

just plug in n= infinity :3

OpenStudy (anonymous):

I think the clue lies in what is the definition of limit. The chicken though that since summation of any numbers of 0 is 0, the sum as n tends to infinity must be 0, right?

OpenStudy (anonymous):

If you prove \[ \sum_{i=1}^nc=c\times n \]It becomes trivial.

OpenStudy (anonymous):

I do not agree with the chicken because the summation is 0 regardless of the number of zeros added. Limiting values are meant to be like asymptotic values that are approached as the number of zeroes increases and not meant to be a constant value obtained regardless of the value for n.

OpenStudy (anonymous):

I don't think the limit is the tricky part.

OpenStudy (anonymous):

yes - hold on... having connection problems atm... let me fix it. x_x

OpenStudy (anonymous):

The chicken is right, but there is something in her proof that is wrong.

OpenStudy (anonymous):

You don't see what's wrong with her proof? Think of circular reasoning. Also what you gave above is a generalization. The proof used for when c= 0 is the same you will use for a general c. You would use proof by induction. Apparently, you didn't seem to notice the mistake is the reasoning.

OpenStudy (anonymous):

It's circular reasoning because the chicken used what she was trying to prove to prove what she needed to prove.

OpenStudy (anonymous):

No the result wasn't used to prove itself.

OpenStudy (anonymous):

\(p(k)\implies p(k+1)\) is part of induction.

OpenStudy (anonymous):

the one underlined is what we are trying to prove. We cannot assume that it's equal to 0 which is what is assumed here. Okay. That's basically the logic to it.

OpenStudy (anonymous):

You assume \(p(k)\) is true and use it to show \(p(k+1)\) is true.

OpenStudy (anonymous):

Also that's the last step of the induction process. She skipped the middle step. In that case, her proof would be correct.

OpenStudy (anonymous):

What is the middle step?

OpenStudy (anonymous):

The middle step is to assume that the general statement is true.

OpenStudy (anonymous):

General statement? There is only the base case and inductive hypothesis. There isn't any middle step.

OpenStudy (anonymous):

2 and 3 are redundant but okay.

OpenStudy (anonymous):

Struggling with latex?

OpenStudy (anonymous):

The middle step is to assume that the general step is to assume that the general statement is true. The last step is to show that if you assume 2) is true. Then it's also true for the next term. General statement is what we are trying to prove. Which is on the board, i'm not sure if it's okay what the chicken say maybe it's assumed that we should know that the general term is assumed to be true.

OpenStudy (anonymous):

lol, I'm using a different software from chrome web store its: Daum equation editor

OpenStudy (anonymous):

Anyways, I'm not sure there is anymore to really say about this.

OpenStudy (anonymous):

I made a new question, look at it!